OpenStudy (anonymous):

2.Solve x^2 = 14x – 49.

OpenStudy (anonymous):

rewrite as \[x^2-14x+49=0\] and then recognize the left hand side as a perfect square

OpenStudy (anonymous):

solve the quadratic

OpenStudy (anonymous):

but i cant figure out the answer

OpenStudy (anonymous):

use x=(-b+/-sqrt(b^2-4ac))/2a

OpenStudy (anonymous):

\[\begin{array}{*{20}c} {x = \frac{{ - b \pm \sqrt {b^2 - 4ac} }}{{2a}}} & {{\rm{when}}} & {ax^2 + bx + c = 0} \\ \end{array} \]

OpenStudy (anonymous):

can you tell me what your a=? b=? c=?

OpenStudy (anonymous):

He just wants the answer don't give it to him

OpenStudy (anonymous):

Show your work I can easily say the answer is 6 and I'm no more right than you.

OpenStudy (inkyvoyd):

Guys, please report jolly. He has been just giving the answer for the past 4 questions he has answered.

OpenStudy (inkyvoyd):

Check closed questions for proof.

OpenStudy (farmdawgnation):

@jollysailorbold What you are doing is called answer sniping and is against our Code of Conduct. Please don't interrupt someone else who is tutoring and just dive-bomb in wit the answer. I'm deleting your replies in this question, which will revoke your medals and SmartScore points. Let me know if you have any questions.